OpenStudy (lina777):

Will award medal and fan! 2NaOH + CO2 --> Na2CO3 + H20 There are 1.85 mol NaOH and 1.00 mol CO2 Question: How many moles of Na2CO3 can be produced??

OpenStudy (photon336):

Okay first thing you need to do is look up the molar masses of NaOH CO2 and Na2CO3

OpenStudy (photon336):

@lina777

OpenStudy (lina777):

@Photon336 NaOH=39.997 CO2=44.01 Na2CO3=105.9888

OpenStudy (lina777):

Thanks for helping by the way :)

OpenStudy (photon336):

So to do this we need to find out which reactant is the limiting reagent this means what reagent will run out first.

OpenStudy (lina777):

I already found this out, and it is the NaOH :)

OpenStudy (photon336):

let me double check this. take a look at what I do

OpenStudy (photon336):

\[1.85~mol~\cancel\NaOH*(\frac{ CO_{2} }{ 2\cancel\NaOH }) = NEED~0.925~mol~CO_{2}\]

OpenStudy (photon336):

\[1.0~mol~\cancel\CO_{2}*(\frac{ 2NaOH }{ mole~\cancel\CO2_{2} }) = NEED~2~mol~NaOH \]

OpenStudy (photon336):

We multiply the number of moles of each reactant that we HAVE by the molar ratio

OpenStudy (photon336):

now let's put it together. the molar ratio will tell us how much we need to react. we compare this to how much we have NEED 2 moles of NaOH 0.925 moles of CO2 HAVE 1.85 moles of NaOH 1.0 mole of CO2 so clearly we need 2 moles of NaOH but we only have 1.85 moles of NaOH Therefore NaOH is going to run out first and so that's the limiting reagent. so now the next step is once we found out what our limiting reagent is (NaOH) we use this to figure out how much product we have.

OpenStudy (photon336):

\[1.85~\cancel\mol~\cancel\NaOH*(\frac{ mol~Na_{2}CO_{3} }{ 2\cancel\NaOH }) = 0.925~mol~Na_{2}CO_{3}\]

OpenStudy (photon336):

The way I set it up was this: You want to make sure that you multiply by the molar ratio. if you notice the compound you want to find is in the top the numerator while the compound you don't want is in the denominator and they cancel out.

OpenStudy (lina777):

This is great! Just a refresher, what do you mean when you say molar ratio?

OpenStudy (photon336):

sorry it's balanced I just checked

OpenStudy (photon336):

You can't do stoichiometry without balancing the equation.

OpenStudy (photon336):

you see those numbers the 2 in front of NaOH and the 1 in front of CO2?

OpenStudy (photon336):

Well molar ratio means the ratio of the number of moles of each compound in the balanced equation

OpenStudy (lina777):

Oh! Got it. Haha. I know what they are, I just forgot what they were called :)

OpenStudy (photon336):

Yeah no problem.

OpenStudy (lina777):

So, the final answer .925 mol Na2CO3? @Photon336

OpenStudy (photon336):

Yes

OpenStudy (lina777):

Okay! Thank you again!

OpenStudy (photon336):

No problem